# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v2

@@ Abstract
The valley degree of freedom in 2D semiconductors, such as gapped bilayer graphene (BLG) and transition metal dichalcogenides, is a promising carrier of quantum information in the emerging field of valleytronics. While valley dynamics have been extensively studied for moderate band gap 2D semiconductors using optical spectroscopy techniques, very little is known about valley lifetimes in narrow band gap BLG, which is difficult to study using optical techniques. Here, we report single-particle valley relaxation times ($T_{1}$) exceeding several microseconds in electrostatically defined BLG quantum dots (QDs) using a pulse-gating technique. The observed dependence of $T_{1}$ on perpendicular magnetic field can be understood qualitatively and quantitatively by a model in which $T_{1}$ is limited by electron-phonon coupling. We identify the coupling to acoustic phonons via the bond length change and via the deformation potential as the limiting mechanisms.
@@
To create gate-defined QDs in BLG, the electronic wave function needs to be confined by a potential $U(\textbf{r})$ in real space (see Fig. 1(a)) and will be distributed near the $K$ and $K^{\prime}$ points, in k-space. A single-electron or single-hole QD can then be described by the Hamiltonian $H_{\rm QD}=H_{\rm BLG}+H_{\rm Z}+H_{\rm SO}+U({\bf r})$ [25]. Here, $H_{\rm BLG}$ represents the effective $4\times 4$ Hamiltonian of bulk BLG near the $K$- and $K^{\prime}$-points based on the sublattice and layer degrees of freedom and includes the bulk valley Zeeman effect (see Supplementary Material), which will be further modified by the confinement $U({\bf r})$. $H_{\rm Z}$ denotes the spin Zeeman coupling. $H_{\rm SO}$ describes the intrinsic Kane-Mele spin-orbit (SO) coupling, which lifts the zero $B$-field degeneracy of the four single-particle states, leading to the formation of two Kramers’ pairs with opposite spin and valley quantum numbers, (|$\ket{K\uparrow}$,|$\ket{K^{\prime}\downarrow}$) and (|$\ket{K^{\prime}\uparrow}$, |$\ket{K\downarrow}$), separated by the SO gap, $\Delta_{\mathrm{SO}}$. In BLG devices $\Delta_{\mathrm{SO}}$ has typically values in the range of $40-80~{}\,\mu$eV [18, 23, 20, 26, 27]. In Fig. 1(d) we show the single-particle spectrum of a BLG hole QD as function of the out-of-plane magnetic field, $B_{\perp}$. As $B_{\perp}$ couples to both the spin and the valley magnetic moments, we observe linear energy shifts given by $E(B_{\perp})-E(0)=(\pm g_{\mathrm{s}}\pm g_{\mathrm{v}})\mu_{\mathrm{B}}B_{\perp}/2$ [28]. Here, $\mu_{\text{B}}$ is the Bohr magneton, $g_{\mathrm{s}}\approx 2$ is the spin g-factor and the valley g-factor, $g_{\mathrm{v}}$, quantifies the strength of the Berry-curvature induced valley magnetic moment, which can be tuned by the confinement potential of the QD in a range typically between $g_{\text{v}}\approx 10$ and 70 [24, 29, 30]. All this makes the valley degree of freedom in BLG QDs highly tunable, in stark contrast to its behavior in Si, Ge, or SiGe QDs, enabling significant valley polarization at relatively low $B_{\perp}$-fields, as evident by the separation between the |$\ket{K^{\prime}}$ and excited |$\ket{K}$ states shown in Fig. 1(d).
@@
The fabricated device consists of a flake of BLG encapsulated by two crystals of hexagonal boron nitride (hBN) and placed on a graphite flake acting as a back gate. On top of the van-der-Waals heterostructure, split gates (SGs) are used to gap out the BLG underneath, resulting in a narrow n-type conductive channel connecting the source (S) and drain (D) leads (see Fig. 2(a)). To confine single charge carriers, the band edge profile along the channel can be adjusted using two layers of interdigitated finger gates (FGs) [19, 17]. One of the FGs is used as a plunger gate (PG) to tune the QD, locally overcompensating for the channel potential set by the back gate. The width of the PG measures about 70 nm and the separation of the SGs is around 80 nm, setting an upper limit of the QD radius $r$ to around 30 to 40 nm. The DC potential applied to the plunger gate, $V_{\text{PG}}$, allows to control the charge carrier occupation down to the last hole (see Fig. 2(b)). To study transient transport through the QD, an AC potential, $V_{\mathrm{AC}}$, can be applied to the PG via a bias tee (see Fig. 2(a)). To maximize the transient currents and to study the relaxation dynamics of the QD states, the FGs adjacent to the PG (yellow and blue) are used to reduce the tunnel coupling between the QD and the left and right reservoir, $\Gamma_{\mathrm{L}}$ and $\Gamma_{\mathrm{R}}$ [31, 32, 33, 34]. Details on the device fabrication and the experimental setup are given in the Supplementary Material.
@@
Figure 3: (a) Top: Schematic of the three-level pulse scheme applied to the PG which is characterized by the voltages $V_{\text{i}},V_{\text{h}}$, $V_{\text{m}}$ and the times $\tau_{\text{i}},\tau_{\text{h}}$ and $\tau_{\text{m}}$. Bottom: Schematic of the QD states relative to the electrochemical potentials in the leads (see text for details). (b) Current $I$ as a function of $V_{\text{PG}}$ while the pulse sequence in a is applied ($B_{\perp}=0.22$ T, $V_{\mathrm{SD}}=10~{}\mu$V). The valley $T_{1}$ time is derived from the amplitude of |$\ket{K\uparrow}_{\text{m}}$. Dashed curves are Lorentzian fits to the peaks. (c) Relative occupation probability of |$\ket{K\uparrow}$ after the holding pulse, $P_{\ket{K\uparrow}}(\tau_{\text{h}})/P_{\ket{K\uparrow}}(0)$ as a function of the holding time, $\tau_{\text{h}}$. The traces are offset for clarity.
@@
Next, to investigate the relaxation dynamics of the observed single-particle valley ES, we apply a three-level pulse scheme to the PG (see Fig. 3(a)) and measure the tunneling current through the QD [33]. During $\tau_{\mathrm{i}}$, the QD is initialized in the empty state. Subsequently, during $\tau_{\mathrm{h}}$, the GS |$\ket{K^{\prime}\uparrow}$, the spin ES |$\ket{K^{\prime}\downarrow}$ and the two valley ESs |$\ket{K\uparrow}$ and |$\ket{K\downarrow}$ are tuned below the electrochemical potentials of the source and drain leads. After the characteristic tunneling time, on the order of $1/(\Gamma_{\mathrm{L}}+\Gamma_{\mathrm{R}})$ either the GS or one of the three ESs will be occupied by a single charge carrier. A charge carrier in an ES has the chance to relax into an energetically lower lying state by either spin or valley relaxation with a characteristic relaxation time, $T_{1}$. Finally, during $\tau_{\mathrm{m}}$, we perform a valley selective readout, measured by aligning the |$\ket{K}$ states in the bias window. Only charge carriers occupying one of the two |$\ket{K}$ states, which have not relaxed into a |$\ket{K^{\prime}}$, can tunnel out and contribute to the transient current. Fig. 3(b) shows the current, $I$, through the QD as a function of $V_{\mathrm{PG}}$ while applying the pulse sequence depicted in Fig. 3(a). The three peaks labeled |$\ket{K^{\prime}\uparrow}_{\mathrm{i}}$, |$\ket{K^{\prime}\uparrow}_{\mathrm{h}}$ and |$\ket{K^{\prime}\uparrow}_{\mathrm{m}}$ correspond to GS transport during each of the three pulse steps. Furthermore, transient currents via the three ESs, |$\ket{K\uparrow}_{\mathrm{m}}$, |$\ket{K\downarrow}_{\mathrm{m}}$ and |$\ket{K^{\prime}\downarrow}_{\mathrm{m}}$, can be observed during $\tau_{\mathrm{m}}$. The relaxation time $T_{1}$ of |$\ket{K\uparrow}$ into an energetically lower lying state can be probed by measuring the amplitude of the |$\ket{K\uparrow}_{\mathrm{m}}$ peak as a function of the holding time $\tau_{\mathrm{h}}$ [31, 33, 34]. We convert the current $I$ into the number of charge carriers tunneling through the QD per pulse cycle, $\langle n\rangle/\text{pulse}=I(\tau_{\text{i}}+\tau_{\text{h}}+\tau_{\text{m}})/e$. The number of charge carriers $\langle n\rangle_{\ket{K\uparrow}_{\mathrm{m}}}$ tunneling via the ES |$\ket{K\uparrow}_{\mathrm{m}}$, is directly proportional to the probability of |$\ket{K\uparrow}$ remaining occupied after $\tau_{\text{h}}$, $P_{\ket{K\uparrow}}(\tau_{\text{h}})$. The relative occupation probability of |$\ket{K\uparrow}$ as a function of $\tau_{\mathrm{h}}$ decays exponentially with the characteristic relaxation time, $T_{1}$, $\langle n\rangle_{\ket{K\uparrow}_{\mathrm{m}}}(\tau_{\mathrm{h}})/\langle n\rangle_{\ket{K\uparrow}_{\mathrm{m}}}(0)=P_{\ket{K\uparrow}}(\tau_{\text{h}})/P_{\ket{K\uparrow}}(0)=e^{-\tau_{\mathrm{h}}/T_{1}}$ [31, 33].
@@
Fig. 3(c) depicts $P_{\ket{K\uparrow}}(\tau_{\text{h}})/P_{\ket{K\uparrow}}(0)$ as function of $\tau_{\mathrm{h}}$ for three different out-of-plane magnetic fields. The datasets show an exponential decay of the occupation probability as a function of $\tau_{\mathrm{h}}$. An exponential fit (solid line), yields for example $T_{1}=4.0~{}\mu$s at $B_{\perp}=0.175$ T. $T_{1}$ decreases with increasing $B_{\perp}$ and reaches a value of 845 ns at $B_{\perp}=0.45$ T. A single charge carrier occupying |$\ket{K\uparrow}$ may relax into a lower lying state either by pure valley relaxation (|$\ket{K\uparrow}\rightarrow\ket{K^{\prime}\uparrow}$) or by additionally flipping the spin (|$\ket{K\uparrow}\rightarrow\ket{K^{\prime}\downarrow}$). Relaxation processes requiring a single valley flip are expected to be faster than processes that require both a spin and valley flip. This is supported by recently reported spin relaxation times between hundreds of microseconds to up to 50 ms, for energy splittings, $\Delta E_{\mathrm{S}}>200$ $\mu$eV [33, 34]. For the pure spin relaxation between |$\ket{K^{\prime}\downarrow}$ and the GS, no relaxation could be observed over the whole range of investigated $\tau_{\mathrm{h}}$ and $B_{\perp}$ (the amplitude of |$\ket{K^{\prime}\downarrow}_{\mathrm{m}}$ is constant as a function of $\tau_{\mathrm{h}}$, see Supplementary Fig. S1). Hence, we conclude that $T_{1}$ extracted from Fig. 3(c) must be limited by the valley relaxation time. In Fig. 4, we plot the valley relaxation time $T_{1}$ extracted from exponential fits, as exemplarily shown in Fig. 3(c) as a function of $B_{\perp}$ and as a function of the energy splitting $\Delta E_{\mathrm{v}}$. When decreasing $B_{\perp}$ from 0.7 T to about 0.15 T, $T_{1}$ increases from below 0.5 $\mu$s to about 7 $\mu$s, while at even lower $B_{\perp}$, the relaxation rate decreases again to $T_{1}\sim$$\mu$s at 80 mT (gray).
@@
We only take into account the emission of phonons (with energy $v_{\mathrm{\lambda}}q$) as the thermal energy is significantly smaller than the valley splitting. To quantify the electron-phonon coupling strength, we perform the least square fit to the experimental data using $g_{1}$ and $g_{2}$ as free fit parameters. Our model is in good qualitative and quantitative agreement with the data taken above $B_{\perp}=0.1$ T, where increasing $B_{\perp}$ results in decreasing $T_{1}$, while it cannot explain the decrease in $T_{1}$ observed for $B_{\perp}<0.1~{}$T, suggesting that other mechanisms dominate in this regime. We speculate that the discrepancy between our model and the data in this regime may be due to a hot spot [40], thermal broadening induced charge noise [41, 42] or $1/f$ charge noise, which is expected to become dominant in the small energy range, as discussed in more detail in ref. [25]. Consequently, we have restricted the fit to the data for $B_{\perp}\geq 0.1$ T, which yields coupling parameters of $g_{1}=50~{}$eV, $g_{2}=5.4~{}$eV. It is noteworthy that both parameters are in good agreement with literature, which includes values in the range of $20$ to $50$ eV for $g_{1}$ [43, 44, 45, 46, 47] and values in the range of $1.5$ to $5$ eV for $g_{2}$ [47, 36], where the wide range of values is partly due to the dependence of the deformation potential on screening and doping [43, 44]. The black solid line in Fig. 4 corresponds to the contributions from both, the deformation potential coupling and from bond length change coupling, while the dashed lines show the individual contributions, respectively (see labels). In the calculation, the intervalley coupling, which is mainly responsible for the absolute values of $T_{1}$ but does not enter the functional $B_{\perp}$ dependence, was set to $\Delta_{KK^{\prime}}=50~{}\mu$eV.

### v3

# Phonon-limited valley lifetimes in single-particle bilayer graphene quantum dots

@@ Abstract
The valley degree of freedom in 2D semiconductors, such as gapped bilayer graphene (BLG) and transition metal dichalcogenides, is a promising carrier of quantum information in the emerging field of valleytronics. While valley dynamics have been extensively studied for moderate band gap 2D semiconductors using optical spectroscopy techniques, very little is known about valley lifetimes in narrow band gap BLG, which is difficult to study using optical techniques. Here, we report single-particle valley relaxation times ($T_{1}$) exceeding several microseconds in electrostatically defined BLG quantum dots using a pulse-gating technique. The observed dependence of $T_{1}$ on perpendicular magnetic field can be understood qualitatively and quantitatively by a model in which $T_{1}$ is limited by electron-phonon coupling. We identify the coupling to acoustic phonons via the bond length change and via the deformation potential as the limiting mechanisms.
@@
To create gate-defined QDs in BLG, the electronic wave function needs to be confined by a potential $U(\textbf{r})$ in real space (see Fig. 1(a)) and will be distributed near the $K$ and $K^{\prime}$ points, in k-space. A single-electron or single-hole QD can then be described by the Hamiltonian $H_{\rm QD}=H_{\rm BLG}+H_{\rm Z}+H_{\rm SO}+U({\bf r})$ [25]. Here, $H_{\rm BLG}$ represents the effective $4\times 4$ Hamiltonian of bulk BLG near the $K$- and $K^{\prime}$-points based on the sublattice and layer degrees of freedom and includes the bulk valley Zeeman effect (see Appendix A), which will be further modified by the confinement $U({\bf r})$. $H_{\rm Z}$ denotes the spin Zeeman coupling. $H_{\rm SO}$ describes the intrinsic Kane-Mele spin-orbit (SO) coupling, which lifts the zero $B$-field degeneracy of the four single-particle states, leading to the formation of two Kramers’ pairs with opposite spin and valley quantum numbers, (|$\ket{K\uparrow}$,|$\ket{K^{\prime}\downarrow}$) and (|$\ket{K^{\prime}\uparrow}$, |$\ket{K\downarrow}$), separated by the SO gap, $\Delta_{\mathrm{SO}}$. In BLG devices $\Delta_{\mathrm{SO}}$ has typically values in the range of $40-80~{}\,\mu$eV [18, 23, 20, 26, 27]. In Fig. 1(d) we show the single-particle spectrum of a BLG hole QD as function of the out-of-plane magnetic field, $B_{\perp}$. As $B_{\perp}$ couples to both the spin and the valley magnetic moments, we observe linear energy shifts given by $E(B_{\perp})-E(0)=(\pm g_{\mathrm{s}}\pm g_{\mathrm{v}})\mu_{\mathrm{B}}B_{\perp}/2$ [28]. Here, $\mu_{\text{B}}$ is the Bohr magneton, $g_{\mathrm{s}}\approx 2$ is the spin g-factor and the valley g-factor, $g_{\mathrm{v}}$, quantifies the strength of the Berry-curvature induced valley magnetic moment, which can be tuned by the confinement potential of the QD in a range typically between $g_{\text{v}}\approx 10$ and 70 [24, 29, 30]. All this makes the valley degree of freedom in BLG QDs highly tunable, in stark contrast to its behavior in Si, Ge, or SiGe QDs, enabling significant valley polarization at relatively low $B_{\perp}$-fields, as evident by the separation between the |$\ket{K^{\prime}}$ and excited |$\ket{K}$ states shown in Fig. 1(d).
@@
The fabricated device consists of a flake of BLG encapsulated by two crystals of hexagonal boron nitride (hBN) and placed on a graphite flake acting as a back gate. On top of the van-der-Waals heterostructure, split gates (SGs) are used to gap out the BLG underneath, resulting in a narrow n-type conductive channel connecting the source (S) and drain (D) leads (see Fig. 2(a)). To confine single charge carriers, the band edge profile along the channel can be adjusted using two layers of interdigitated finger gates (FGs) [19, 17]. One of the FGs is used as a plunger gate (PG) to tune the QD, locally overcompensating for the channel potential set by the back gate. The width of the PG measures about 70 nm and the separation of the SGs is around 80 nm, setting an upper limit of the QD radius $r$ to around 30 to 40 nm. The DC potential applied to the plunger gate, $V_{\text{PG}}$, allows to control the charge carrier occupation down to the last hole (see Fig. 2(b)). To study transient transport through the QD, an AC potential, $V_{\mathrm{AC}}$, can be applied to the PG via a bias tee (see Fig. 2(a)). To maximize the transient currents and to study the relaxation dynamics of the QD states, the FGs adjacent to the PG (yellow and blue) are used to reduce the tunnel coupling between the QD and the left and right reservoir, $\Gamma_{\mathrm{L}}$ and $\Gamma_{\mathrm{R}}$ [31, 32, 33, 34]. Details on the device fabrication and the experimental setup are given in the Appendix B and C.
@@
Figure 3: (a) Top: Schematic of the three-level pulse scheme applied to the PG which is characterized by the voltages $V_{\text{i}},V_{\text{h}}$, $V_{\text{m}}$ and the times $\tau_{\text{i}},\tau_{\text{h}}$ and $\tau_{\text{m}}$. Bottom: Schematic of the QD states relative to the electrochemical potentials in the leads (see text for details). (b) Current $I$ as a function of $V_{\text{PG}}$ while the pulse sequence in a is applied ($B_{\perp}=0.22$ T, $V_{\mathrm{SD}}=10~{}\mu$V). The valley $T_{1}$ time is derived from the amplitude of |$\ket{K\uparrow}_{\text{m}}$. Dashed curves are Lorentzian fits to the peaks. (c) Relative occupation probability of |$\ket{K\uparrow}$ after the holding pulse, $P_{\ket{K}}(\tau_{\text{h}})/P_{\ket{K}}(0)$ as a function of the holding time, $\tau_{\text{h}}$. The traces are offset for clarity.
@@
Next, to investigate the relaxation dynamics of the observed single-particle valley ES, we apply a three-level pulse scheme to the PG (see Fig. 3(a)) and measure the tunneling current through the QD [33]. During $\tau_{\mathrm{i}}$, the QD is initialized in the empty state. Subsequently, during $\tau_{\mathrm{h}}$, the GS |$\ket{K^{\prime}\uparrow}$, the spin ES |$\ket{K^{\prime}\downarrow}$ and the two valley ESs |$\ket{K\uparrow}$ and |$\ket{K\downarrow}$ are tuned below the electrochemical potentials of the source and drain leads. After the characteristic tunneling time, on the order of $1/(\Gamma_{\mathrm{L}}+\Gamma_{\mathrm{R}})$ either the GS or one of the three ESs will be occupied by a single charge carrier. A charge carrier in an ES has the chance to relax into an energetically lower lying state by either spin or valley relaxation with a characteristic relaxation time, $T_{1}$. Finally, during $\tau_{\mathrm{m}}$, we perform a valley selective readout, measured by aligning the |$\ket{K}$ states in the bias window. Only charge carriers occupying one of the two |$\ket{K}$ states, which have not relaxed into a |$\ket{K^{\prime}}$, can tunnel out and contribute to the transient current. Fig. 3(b) shows the current, $I$, through the QD as a function of $V_{\mathrm{PG}}$ while applying the pulse sequence depicted in Fig. 3(a). The three peaks labeled |$\ket{K^{\prime}\uparrow}_{\mathrm{i}}$, |$\ket{K^{\prime}\uparrow}_{\mathrm{h}}$ and |$\ket{K^{\prime}\uparrow}_{\mathrm{m}}$ correspond to GS transport during each of the three pulse steps. Furthermore, transient currents via the three ESs, |$\ket{K\uparrow}_{\mathrm{m}}$, |$\ket{K\downarrow}_{\mathrm{m}}$ and |$\ket{K^{\prime}\downarrow}_{\mathrm{m}}$, can be observed during $\tau_{\mathrm{m}}$. The relaxation time $T_{1}$ of the |$\ket{K}$ states into an energetically lower lying state can be probed by extracting the maximum value of the combined |$\ket{K\uparrow}_{\mathrm{m}}$ and |$\ket{K\downarrow}_{\mathrm{m}}$ peaks, which at a fixed value of $B_{\perp}$ is at a constant energy difference (see Fig. 1(d)), as a function of the holding time $\tau_{\mathrm{h}}$ (for more information on the peak analysis see Appendix D) [31, 33, 34]. We convert the current $I$ into the number of charge carriers tunneling through the QD per pulse cycle, $\langle n\rangle/\text{pulse}=I(\tau_{\text{i}}+\tau_{\text{h}}+\tau_{\text{m}})/e$. The number of charge carriers $\langle n\rangle_{\ket{K}_{\mathrm{m}}}$ tunneling via the excited |$\ket{K}$ states, is directly proportional to the probability of |$\ket{K}$ remaining occupied after $\tau_{\text{h}}$, $P_{\ket{K}}(\tau_{\text{h}})$. The relative occupation probability of |$\ket{K}_{\mathrm{m}}$ as a function of $\tau_{\mathrm{h}}$ decays exponentially with the characteristic relaxation time, $T_{1}$, $\langle n\rangle_{\ket{K}_{\mathrm{m}}}(\tau_{\mathrm{h}})/\langle n\rangle_{\ket{K}_{\mathrm{m}}}(0)=P_{\ket{K}}(\tau_{\text{h}})/P_{\ket{K}}(0)=e^{-\tau_{\mathrm{h}}/T_{1}}$ [31, 33].
@@
Fig. 3(c) depicts $P_{\ket{K}}(\tau_{\text{h}})/P_{\ket{K}}(0)$ as function of $\tau_{\mathrm{h}}$ for three different out-of-plane magnetic fields. The datasets show an exponential decay of the occupation probability as a function of $\tau_{\mathrm{h}}$. An exponential fit (solid line), yields for example $T_{1}=4.0~{}\mu$s at $B_{\perp}=0.175$ T. $T_{1}$ decreases with increasing $B_{\perp}$ and reaches a value of 845 ns at $B_{\perp}=0.45$ T. A single charge carrier occupying |$\ket{K\uparrow}$ or |$\ket{K\downarrow}$ may relax into a lower lying state either by pure valley relaxation (|$\ket{K\uparrow}\rightarrow\ket{K^{\prime}\uparrow}$ and |$\ket{K\downarrow}\rightarrow\ket{K^{\prime}\downarrow}$) or by additionally flipping the spin (|$\ket{K\uparrow}\rightarrow\ket{K^{\prime}\downarrow}$ and |$\ket{K\downarrow}\rightarrow\ket{K^{\prime}\uparrow}$). Relaxation processes requiring a single valley flip are expected to be faster than processes that require both a spin and valley flip. This is supported by recently reported spin relaxation times between hundreds of microseconds to up to 50 ms, for energy splittings, $\Delta E_{\mathrm{S}}>200$ $\mu$eV [33, 34]. For the pure spin relaxation between |$\ket{K^{\prime}\downarrow}$ and the GS, no relaxation could be observed over the whole range of investigated $\tau_{\mathrm{h}}$ and $B_{\perp}$ (the amplitude of |$\ket{K^{\prime}\downarrow}_{\mathrm{m}}$ is constant as a function of $\tau_{\mathrm{h}}$, see Appendix D). Hence, we conclude that $T_{1}$ extracted from Fig. 3(c) must be limited by the valley relaxation time.
In Fig. 4, we plot the valley relaxation time $T_{1}$ extracted from exponential fits, as exemplarily shown in Fig. 3(c) as a function of $B_{\perp}$ and as a function of the energy splitting $\Delta E_{\mathrm{v}}$. When decreasing $B_{\perp}$ from 0.7 T to about 0.15 T, $T_{1}$ increases from below 0.5 $\mu$s to about 7 $\mu$s, while at even lower $B_{\perp}$, the relaxation rate decreases again to $T_{1}\sim$$\mu$s at 80 mT (gray).
@@
We only take into account the emission of phonons (with energy $v_{\mathrm{\lambda}}q$) as the thermal energy is significantly smaller than the valley splitting. To quantify the electron-phonon coupling strength, we perform the least square fit to the experimental data using $g_{1}$ and $g_{2}$ as free fit parameters. Our model is in good qualitative and quantitative agreement with the data taken above $B_{\perp}=0.1$ T, where increasing $B_{\perp}$ results in decreasing $T_{1}$, while it cannot explain the decrease in $T_{1}$ observed for $B_{\perp}<0.1~{}$T, suggesting that other mechanisms dominate in this regime. We speculate that the discrepancy between our model and the data in the regime $B_{\perp}<0.1~{}$T may be due to a hot spot [40], thermal broadening induced charge noise [41, 42] or $1/f$ charge noise, which is expected to become dominant in the small energy range, as discussed in more detail in ref. [25]. Consequently, we have restricted the fit to the data for $B_{\perp}\geq 0.1$ T, which yields coupling parameters of $g_{1}=50~{}$eV, $g_{2}=5.4~{}$eV. It is noteworthy that both parameters are in good agreement with literature, which includes values in the range of $20$ to $50$ eV for $g_{1}$ [43, 44, 45, 46, 47] and values in the range of $1.5$ to $5$ eV for $g_{2}$ [47, 36], where the wide range of values is partly due to the dependence of the deformation potential on screening and doping [43, 44]. The black solid line in Fig. 4 corresponds to the contributions from both, the deformation potential coupling and from bond length change coupling, while the dashed lines show the individual contributions, respectively (see labels). In the calculation, the intervalley coupling, which is mainly responsible for the absolute values of $T_{1}$ but does not enter the functional $B_{\perp}$ dependence, was set to $\Delta_{KK^{\prime}}=50~{}\mu$eV.